If the sides of a square are decreased by 2cm, the area is decreased by 36cm2. What were the dimensions of the original square?
Answer by mananth(16946)

let side of square be x
Area = x^2
it is decreased by 2
so new length = x-2
area = (x-2)^2
..
The area decreases by 36 m^2
x^2-(x-2)^2=36
x^2-x^2+2x-4=36
2x-4 = 36
2x=40
x= 20 m
...
side = 20m
...
m.ananth@hotmail.ca